The sum of two of the numbers is six. The sum of the squares of the two numbers is twenty. Find the two numbers.
Okay, so far I was thinking the first equation would look like x+x=6, and the second equation would look like (x+x)^2=20. Is this correct? Thank you for your time!

Answer by Alan3354(69443)

It has to be 2 different numbers.
x + x = 6 --> x = 3
------------
Then, you have (x + x)^2 = 20. That's not what it says.
------------
x = 1 number
y = the other number
x + y = 6
x^2 + y^2 = 20 not (x+y)^2 = 20
-------------
Email me at Moral Loophole@aol.com for further assistance, if needed.